Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
-3x + 6y = -3
-6x + 4y = -10

Answer by Alan3354(69443)

-3x + 6y = -3
-6x + 4y = -10
---------------
divide the 2nd eqn by 2, then subtract.
Solve for y.
Sub for y to solve for x.
Show your work.